### v1

# K-COMP: Retrieval-Augmented Medical Domain Question Answering With Knowledge-Injected Compressor

Abstract
Retrieval-augmented question answering (QA) integrates external information, and thereby increases the QA accuracy of reader models that lack domain knowledge. However, documents retrieved for closed domains require high expertise, so the reader model may have difficulty fully comprehending the text. Moreover, the retrieved documents contain thousands of tokens, some unrelated to the question. As a result, the documents include some inaccurate information, which could lead the reader model to mistrust the passages and could result in hallucinations. To solve these problems, we propose K-comp (Knowledge-injected COMPressor) which provides the knowledge required to answer correctly. The compressor automatically generates the requisite prior knowledge to facilitate the answering process prior to the compression of retrieved passages. Subsequently, the passages are compressed autoregressively, with the generated knowledge being integrated into the compression process. This process ensures alignment between the question intent and the compressed context. By augmenting this prior knowledge and concise context, the reader models are guided toward relevant answers and trust the context.
K-COMP: Retrieval-Augmented Medical Domain Question Answering With Knowledge-Injected Compressor
Jeonghun Cho1  Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence 2Department of Computer Science and Engineering Pohang University of Science and Technology, Republic of Korea {jeonghuncho, gblee}@postech.ac.kr
1 Introduction
Retrieval-augmented question answering (QA) is a task where passages related to a question are appended to the prompt such that a reader model can reference them and the infer correct answer (Ahmad et al., 2019; Guo et al., 2021).
Figure 1: K-comp helps the reader model infer accurate responses by using domain knowledge and compressed context aligned with the question.
However, several limitations impede retrieval-augmented approaches in closed domains with large language models (LLMs) as readers. First, the documents retrieved for closed domains require domain expertise, so the reader may not trust the whole text. When faced with unfamiliar input, the model exhibits an availability bias towards commonly known knowledge, making it more willing to believe in information they can easily recall (Jin et al., 2024). Also, retrieved passages contain thousands of tokens and are sometimes unrelated to the question, which can cause the language model to distrust the passages, perceive them as irrelevant noise, and generate answers that do not consider them. These problems lead to hallucinations (Ji et al., 2023a), which result in the model generating inaccurate answers or inferring plausible but false responses. Lastly, LLMs are sensitive to the order of retrieved documents and the prompting method. Specifically, LLMs can have difficulty finding the necessary information within lengthy input prompts, especially when key information or correct answer clues are located in the middle of the prompt (Liu et al., 2024; Xu et al., 2024b).
To tackle these issues, we propose K-comp (knowledge-injected compressor). We aim to use an autoregressive LLM as a compressor with the domain knowledge needed to answer the question, and increase the alignment of the retrieved passages with the question intent. Additionally, when the compressor is trained domain-related terms and information, it becomes able to recognize the entities that occur in the question and provide descriptions for them. This process is significant for closed domains that require substantial prior knowledge. For retrieval augmentation, we use a large amount of text from domain-specific sources, including Wikipedia. We exploit the advantages of domain relevance by efficiently reusing it when annotating prior knowledge, not just for retrieval. Furthermore, we use a causal masking objective (Aghajanyan et al., 2022) during the training phase to inject domain knowledge into the compressor.
In summary, our contributions are as follows:
  • We propose a novel approach to generate knowledge-injected summaries adapted for the medical domain. We incorporate causal masking to inject knowledge into the compressor without modifying its structure. This approach ensures that the summary is aligned with the question.
  • Even without domain knowledge in the reader model, K-comp provides the medical jargon description to answer the question, thereby enabling LLMs with diverse backgrounds to handle medical questions more accurately.
  • Experiments on three medical datasets show that K-comp improves performance over other query-based prompt compression methodologies and standard retrieval-augmented generation (RAG) without compression.
  • K-comp has been shown to be effective when applied to previously unseen data, thereby presenting evidence that our method provides additional novel contributions in data-scarce closed domain environments.
2 Related Work
Text Infilling.
Models such as BERT (Devlin et al., 2019), SpanBERT (Joshi et al., 2020), T5 (Raffel et al., 2020), and BART (Lewis et al., 2020), are pre-trained using masked language modeling within a bidirectional encoder architecture. They have shown strong performance in infilling short and contiguous masked token spans. However, the bidirectional attention mechanism typically restricts the fillable span length to dimensions significantly shorter than a sentence.
In contrast, decoder-only models such as GLM (Du et al., 2022), CM3 (Aghajanyan et al., 2022), and InCoder (Fried et al., 2023) operate by left-to-right generation. They can accommodate variable infill span lengths. Causal masking (Aghajanyan et al., 2022) or fill-in-the-middle (Bavarian et al., 2022) methods predict masked spans from the posterior context. These methods have their generative capabilities, which increase the length of infill spans. They can also exploit the advantages of considering contextual relationships that surround the masked span. The proposed method has the capability to fill the span by considering bidirectional context, as well as align the generated summary with the question by regressively encoding the infilled span.
Prompt Compression.
Several studies have demonstrated that prompt augmentations effectively enhance the performance of LLMs across various tasks (Liu et al., 2023a; Ram et al., 2023; Ryu et al., 2023; Wang et al., 2024c; Long et al., 2023; Yagnik et al., 2024). Yet, the relevance and reliability of the augmented passages are significant challenges in prompt augmentations. In order to address this issue, recent studies have attempted to directly extract content from ambiguous and lengthy passages. Kim et al. (2024) eliminates irrelevant information while maximizing the extraction of accurate information, whereas Yang et al. (2023) leverages the black-box LLMs by applying a reward-based method during compressor training to generate summaries. RECOMP (Xu et al., 2024a) selects and augments the summary with the highest end-task performance by using prompts in which non-essential summaries are set to empty strings if necessary. LLMLingua (Jiang et al., 2023a) dynamically assigns different compression rates to various components within the prompt. In contrast, K-comp focuses on the keywords needed to answer the question, emphasizing the alignment between the compressed context and the question.
3 Causal Knowledge Injection
Causal models trained using autoregressive language modeling rely exclusively on the context to the left of the generated tokens to predict subsequent tokens (Brown et al., 2020). This attribute confers an advantage in causally generating entire documents, such as text generation. However, these models show limited proficiency in tasks that require an understanding of post-positional relationships for span infilling. Conversely, masked language models excel at predicting masked spans by referencing attention scores from tokens located both anteriorly and posteriorly. Nonetheless, their training objective is limited to decoding only short segments of the passages (Devlin et al., 2019; Joshi et al., 2020).
We are inspired by causal masking (Aghajanyan et al., 2022) that combines the advantages of both objectives. We focus on the masked medical entities within the question (prior context) and aim to predict them by considering the retrieved snippets (subsequent context). Subsequently, by auto-regressively compressing the retrieved snippets, we can effectively leverage both advantages.
4 Methods
In this section, we report our proposed approach for knowledge-injected compression and retrieval augmentation. To retrieve passages similar to a question, we construct a retrieval pipeline composed of a large corpus (§4.1). Next, we explain the data processing steps for training (§4.2). Finally, we detail the training scheme for K-comp with the proposed objective and explain the inference phase for retrieval augmentation (§4.3). Figure 1 shows an overview of the prompts that K-comp consists of.
4.1 Retrieval Framework
Closed domain tasks have not been as thoroughly explored as open domain tasks, which have achieved notable performance enhancements using Wikipedia as a retrieval corpus (Karpukhin et al., 2020). In contrast to open domains, the challenge in closed domains is that unified corpora have not been established. Research endeavors, such as Xiong et al. (2024); Wang et al. (2024b), are currently underway to address this gap. To ensure coverage of both general and domain knowledge, we adopt the MedCorp corpus (Xiong et al., 2024) as our retrieval corpus. It combines Wikipedia, PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/, StatPearls [^2] https://www.statpearls.com/, and textbooks (Jin et al., 2021). As our retriever, we employ embedding-based $k$-NN search (Johnson et al., 2019) to mitigate bottlenecks and efficiently execute similarity searches on our large-scale corpus comprising four distinct text corpora [^3] We use Nomic Embed (Nussbaum et al., 2024)..
4.2 Ground-Truth Data
Entity-Description.
We rely on off-the-shelf tools to perform named-entity recognition [^4] We use ScispaCy (Neumann et al., 2019) package., which identifies biomedical entities =$\mathcal{E}=\{e_{i}\}$ in each question for masking. Retrieval corpus $\mathbb{C}$ is constituted of title and text pairs, with the first sentence of each text assumed to be a short description of the title (Xu et al., 2023). Subsequently, the pairs of titles and short descriptions are matched with the entities and their corresponding knowledge $d_{i}$. We assume that the questions in the training dataset contain at least one entity. In the absence of an entity in a given question, the data are excluded. Similarly, instances lacking a corresponding title in the retrieval corpus are also filtered out of the training dataset (Table 15).
However, in the test data, K-comp unveils a novel contribution by automatically generating domain-specific entity descriptions during inference even when no annotate exists for the entity in question. This obviates the need for costly and unnecessary tasks, such as searching for medical terms or finding definitions within the corpus.
Summary.
To synthesize gold summaries $\mathcal{S}$, GPT-4o-mini [^5] We use gpt-4o-mini-2024-07-18 (OpenAI, 2024). compresses the passages by considering $\{\mathcal{P},\mathcal{E}\}$ input pairs, and the number of passages used for synthesis is set to five, i.e., $|\mathcal{P}|=5$. Notably, we explicitly prohibit the inclusion of the question in the summary synthesis process. This is because incorporating the question into the input prompt for generating the summary may result in a focus shift from the generation of keyword-focused summaries to the formulation of a summary that is aimed at answering the question. Detailed instructions for the summary synthesis are provided in Table 8.
4.3 K-comp
Preliminary.
[…]
$q_{m}=\text{What are the {<ent>} of {<ent>}?}\end{aligned}$
$d_{1}=\text{symptom: {\{\{description\}\}<eod>}}\end{aligned}$
$d_{2}=\text{Down syndrome: {\{\{description\}\}<eod>}}\end{aligned}$
By concatenating $q_{m}$ and $\mathcal{P}$ in the correct sequence, the masked spans can be predicted based on the preceding and subsequent context. We define the dataset for the compressor as $(q_{m}\oplus\mathcal{P},\mathcal{S},\mathcal{E},\mathcal{D})$, where $\mathcal{D}=\{d_{i}\}$ and $\mathcal{S}$ is a gold summary.
Training.
Given an input query $q_{m}$ and the set of retrieved passages $\mathcal{P}=\{p_{1},p_{2},...,p_{5}\}$, K-comp aims to train a causal model $f(q_{m}\oplus\mathcal{P})$ to generate $\mathcal{E}$, $\mathcal{D}$, and then $\mathcal{S}$ auto-regressively.
The compressor is trained to encode $q_{m}\oplus\mathcal{P}$ to generate <ent> tokens and their corresponding descriptions:
$P_{\theta}(\mathcal{E},\mathcal{D}|q_{m}\oplus\mathcal{P})\\ =\prod_{i}\left(\prod_{\alpha,\beta}P_{\theta}(e_{i}^{\alpha},d_{i}^{\beta}|e_{i}^{<\alpha},d_{i}^{<\beta},q_{m}\oplus\mathcal{P})\right)$ (1)
where $\theta$ represents the parameters of K-comp.
This approach facilitates the incorporation of descriptions into the prompt for the reader model and ensures that the generated entities and their descriptions are regressively encoded. As a result, a summary is generated in a causal manner with attention on the entities within the question and their related knowledge, thereby composing a summary centered on these domain entities.
$P_{\theta}(s|\mathcal{E},\mathcal{D},q_{m}\oplus\mathcal{P})\\ =\prod_{\gamma}P_{\theta}(s^{\gamma}|s^{<\gamma},\mathcal{E},\mathcal{D},q_{m}\oplus\mathcal{P})$ (2)
We fine-tuned the compressor using the standard next token prediction with cross-entropy loss:
$P_{\theta}(\mathcal{E},\mathcal{D},s|q_{m}\oplus\mathcal{P})\\ =P_{\theta}(\mathcal{E},\mathcal{D}|q_{m}\oplus\mathcal{P})\times P_{\theta}(s|\mathcal{E},\mathcal{D},q_{m}\oplus\mathcal{P})$ (3)
$\therefore L(\theta)=-\mathbb{E}(\log P_{\theta}(\mathcal{E},\mathcal{D},s\mid q_{m}\oplus\mathcal{P}))$
Inference.
At inference time, documents are retrieved in advance to construct the compressor input batch $\{\mathcal{P},q_{m}\}$. This enables the sequential autoregressive generation of the entities and descriptions from the question until the <eod> token is produced. The overall context, including entities and descriptions, is then considered, and a summary that aligns more closely with the question is generated. This process ultimately constructs the input prompt for the reader model, ensuring a reliable response to the question.
For all datasets, we use a 0-shot setting in our experiments. The prompt examples for the reader model can be found in Table 11.
General-purpose LLMs Medical-purpose LLMs
Llama-3-8B Llama-3-70B Mixtral-8x7B GPT-4o 7 MedAlpaca-13B Meditron-70B
BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval
MedQuAD
Without compressor
Top-1 document 75.65 53.44 79.44 57.47 58.68 44.30 84.45 60.24 75.05 36.89 76.38 52.74
Top-5 documents 76.20 52.45 79.95 55.25 64.17 47.47 83.08 64.98 17.57 10.82 73.14 51.88
With compressor
RECOMP 73.27 55.74 79.75 63.57 71.58 58.78 85.20 60.92 82.52 39.15 76.41 56.14
LLMLingua 76.45 54.65 78.78 57.00 59.79 45.34 84.46 55.04 83.10 44.17 74.74 54.18
FineTune 74.27 54.89 79.97 59.67 71.44 56.26 83.47 62.40 82.25 39.40 76.14 55.97
K-comp 83.79 64.05 82.06 62.91 80.64 63.07 85.49 63.16 85.75 49.39 78.88 58.95
MASH-QA
Without compressor
Top-1 document 77.87 54.08 82.65 56.48 72.63 51.88 83.39 60.66 76.95 43.41 80.45 55.31
Top-5 documents 77.59 52.94 82.07 57.20 73.09 53.18 83.12 64.81 31.91 20.07 79.31 55.85
With compressor
RECOMP 80.20 58.80 82.27 60.50 82.12 62.35 83.82 62.22 78.51 34.83 81.54 60.09
LLMLingua 79.74 54.37 83.25 56.83 76.71 58.32 83.37 56.73 80.41 48.45 80.78 58.55
FineTune 77.30 55.21 82.90 60.66 78.74 60.62 83.88 62.82 79.96 44.02 81.70 60.06
K-comp 83.45 61.45 83.64 61.40 81.49 62.91 83.92 64.06 81.95 51.76 82.67 61.85
BioASQ
Without compressor
Top-1 document 82.53 56.03 81.40 59.58 82.23 62.07 87.13 62.47 76.86 40.06 83.43 57.55
Top-5 documents 81.44 61.33 80.91 61.09 83.54 65.53 86.81 66.15 20.93 12.41 83.63 59.52
With compressor
RECOMP 83.89 59.80 85.45 59.65 86.85 62.82 87.87 60.62 78.07 30.68 84.77 57.58
LLMLingua 81.70 53.62 86.53 61.53 82.54 62.24 86.85 51.11 81.19 42.97 81.51 55.26
FineTune 82.26 60.89 86.17 62.86 85.48 64.01 87.97 63.99 78.67 41.49 85.05 61.20
K-comp 86.61 62.28 87.03 62.89 86.52 66.31 87.77 64.38 84.88 45.47 86.70 62.90
Table 1: Main results. We report automatic evaluations for retrieval-augmented QA with and without compressors.
5 Experiments
In this section, we evaluate K-comp trained by causal knowledge injection and the retrieval-augmented QA task. We report the datasets and settings used in the experiments (§5.1) and discuss the main results (§5.2).
5.1 Settings
Models.
We fine-tuned Gemma-2B (Team et al., 2024) with our knowledge injection objective. Further details regarding the models and implementation can be found in Appendix A.
Datasets.
To reduce potential biases from fine-tuned medical LLMs (Han et al., 2023; Chen et al., 2023), we conduct experiments using the medical QA datasets MedQuAD (Ben Abacha and Demner-Fushman, 2019), MASH-QA (Zhu et al., 2020), and BioASQ (Krithara et al., 2023), which were not directly used for training biomedical models. Although MASH-QA and BioASQ provide gold passages containing answers, our experiments do not utilize these gold passages. Instead, we rely on passages retrieved by our retrieval framework.
Evaluation Metrics.
Since all datasets consist of long-form answers, we use the trained model to evaluate answers. We quantify the relevance of answers by using BertScore (Zhang* et al., 2020), which evaluates the similarity between two sentences by exploiting the contextual embeddings of the encoder. We also use UniEval (Zhong et al., 2022), which is a multi-dimensional evaluation metric that has high correlation and similarity with human judgment. We explicitly assess the factual consistency between generated and gold answers.
5.2 Results
Baselines.
We compare K-comp with standard RAG approach with top-1 and top-5 retrieved passages without applying prompt compression. We also compare with previous state-of-the-art prompt compression methods, including RECOMP (Xu et al., 2024a) and LLMLingua (Jiang et al., 2023a). Specifically, for implementing RECOMP, we use an abstractive compressor fine-tuned on our datasets, and for LLMLingua, we use Llama-2-7B (Touvron et al., 2023) for compression. The prompts for synthesizing the summaries used in RECOMP are based on the paper and can be found in Table 9. Furthermore, the efficacy of causal knowledge injection is evaluated by comparing it to a model that has been fine-tuned (FineTune) using only the standard language modeling objective for summarization. FineTune fine-tuned with Gemma-2B, the same as K-comp.
Overall Performance.
Table 1 shows the main results of K-comp compared to the baselines across various reader LLMs. Overall, compression methods are effective. Chunking snippets for retrieval is inherently imperfect, making the Top-1 and Top-5 passages suboptimal. For MedAlpaca, which has the smallest context window size of 2048 among the reader models, answer accuracy declines significantly with Top-5 passages input due to the limited window size. Consequently, a reprocessing stage, such as compression, is required to improve the quality of chunked text and enable the reader model to reference it appropriately. Among baselines, LLMLingua lags behind other baselines trained in the medical domain due to its query-agnostic compression approach. We also observe different results depending on the model size. Larger models are less dynamic in their response, relying more on their internal knowledge and less on the prompt variations. As can be seen in the case study (§6.3) and Table 14, larger models are capable of providing reasonable responses to questions even when presented with noisy input prompts. In other words, the result demonstrates that not only small models but also large models place greater trust in the description and concise text provided by K-comp than in other baselines.
To emphasize the importance of entity and description, we analyze a scenario where K-comp infers normally, but only appends the summary to the reader prompt (Table 2). $-Prior$ is comparable to the baseline fine-tuned for summarization tasks. Even so, it is clear that providing the reader model with prior knowledge significantly improves the accuracy of the final responses compared to FineTune. An interesting observation is that general-purpose models exhibit a more significant influence of information related to medical jargon compared to medical LLMs. Medical LLMs seem to treat knowledge about entities as noisy input, resulting in conflicts with their internal knowledge. Still, these analyses are confined to the QA accuracy of reader LLMs, as they are affected by changes in the components that make up the prompt. The following sections will discuss the relevance and alignment of the summary.
Llama-3-8B Llama-3-70B MedAlpaca-13B Meditron-70B
BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval
MedQuAD
K-comp 83.79 64.05 82.06 62.91 85.75 49.39 78.88 58.95
$-Prior$ 73.93 54.47 80.06 60.03 84.84 48.23 79.80 59.00
FineTune 74.27 54.89 79.97 59.67 82.25 39.40 76.14 55.97
MASH-QA
K-comp 83.45 61.45 83.64 61.40 81.95 51.76 82.67 61.85
$-Prior$ 78.63 57.80 83.16 60.88 82.18 52.02 81.93 60.42
FineTune 77.30 55.21 82.90 60.66 79.96 44.02 81.70 60.06
BioASQ
K-comp 86.61 62.28 87.03 62.89 84.88 45.47 86.70 62.90
$-Prior$ 80.75 55.06 86.22 62.40 84.53 46.67 85.23 59.73
FineTune 82.26 60.89 86.17 62.86 78.67 41.49 85.05 61.20
Table 2: Ablation studies. $-Prior$ denotes the scenario where K-comp does not provide prior knowledge to the reader LLMs.
Figure 2: Percentage of Recall@$K$ according to the variation of $K$ for the retrieved passages and our compressed contexts, where Top-10 denotes the ten retrieved passages with the highest similarity scores.
6 Analyses
We analyze the results from various perspectives (§6.1, 6.2, 6.3). Finally, we appraise the outcomes focusing on the GPT-4o evaluation to highlight the advantages of K-comp through comparison with previous studies (§6.4).
6.1 Reranking Preference
In addition to evaluating the end-task performance, it is crucial in RAG to ensure that prompts are augmented to be pertinent to the question. Although human evaluation is valuable, it demands significant resources and domain expertise, which are not readily available in our case. Instead, we propose to employ a state-of-the-art reranker [^6] We use BAAI/bge-reranker-large (Xiao et al., 2024). to measure the relevance between the context and the question. For each question $q$, we execute K-comp to generate 10 contexts using a high temperature setting (temperature=1) based on $q_{m}\oplus\mathcal{P}$. Next, we retrieve the top-10 passages related to $q$. Thus, we gather a total of 20 passages to be fed to the reranker. By applying Recall@$K$ to these 20 passages, we observe the $K$ passages that are most similar to $q$, and quantify the proportion of the compressor varied as $K$ varied.
Figure 2 illustrates Recall@$K$ across different values of $K$. Specifically, we achieved Recall@1 scores of 77%, 73%, and 83% on MedQuAD, MASH-QA, and BioASQ, whereas the top-5 retrieved passages achieved 23%, 27%, and 17%. This comparison demonstrates that the reranker strongly prefers our compressed contexts across all three benchmarks.
Settings Top-1 Top-5 K-comp
Input tokens 195 978 198
Inference time 9,798s 20,193s 9,777s
Compression time - - 497s
Total time 9,798s 20,193s 10,274s
Table 3: Inference speed of Llama-3-70B on MASH-QA.
Question: How do x-rays diagnose rheumatoid arthritis (RA)?
Augmented Prompt Answer
FineTune ### Passage
Rheumatoid arthritis (RA) is a chronic systemic disease characterized by inflammation of the synovial tissue, affecting 0.5%-1.0% of the global population, with …
X-rays are used to diagnose RA by detecting joint damage and erosion.
K-comp ### Entity
x-ray:{{description}}
rheumatoid arthritis: {{description}}
### Passage
X-rays are the most common method for assessing the degree of joint destruction in rheumatoid arthritis (RA), revealing osteopenia and joint space narrowing as the disease …
X-rays diagnose rheumatoid arthritis (RA) by revealing changes in the axial skeleton and joint space. As the disease progresses, early RA may not show significant changes on X-rays, but it can indicate osteopenia and joint space narrowing. Advances in imaging techniques like MRI and ultrasound enhance the ability to detect soft tissue changes and joint damage, which is crucial for accurate diagnosis …
Table 4: Case study. We provide the passages used to augment the reader prompt and the answers. Red texts highlight the medical jargon within the question. The complete prompt can be found in Table 14
6.2 Inference Speed
In Table 3, we report the inference time and the number of tokens used in the prompt input as metrics for evaluating efficiency. Specifically, we employed Llama-3-70B as the reader model and measured the GPU runtime on MASH-QA test set. Both the compressor and reader are executed on a single NVIDIA A100 GPU with 80GB memory. Even when considering the time needed for the compressor inference, our method was able to double the throughput compared to prepending the top-5 passages, making it more efficient. Moreover, we note that inference speed is dependent on the implementation and size of the reader model. For instance, models with more parameters will suffer increased latency by increasing the number of input tokens. This phenomenon amplifies the speed advantage of K-comp.
6.3 Case Study
Here, we report how K-comp generates medical knowledge. In Table 4, K-comp is able to address the practical utilization of X-rays in the diagnosis and monitoring of RA, and provides detailed explanations regarding their application, which offers a more transparent rationale for addressing the questions. By contrast, FineTune provides a more general context without focusing on the specific role of X-rays in RA diagnosis. Although it mentions X-rays along with other diagnostic techniques, its focus is on advancements in imaging methods such as MRI. FineTune merely summarises the passages retrieved based on semantic and overall lexical similarities to the question without considering the queried intent. This results in the reader model does not fully trusting the augmented passages, instead perceiving them as irrelevant noise and generating answers not based on the passages. This result can lead to inaccuracies and potential hallucinations.
6.4 GPT-4o Evaluation
We additionally explore the reliability of the context. Given that GPT-4 has been demonstrated to correlate highly with human judgments (Liu et al., 2023b), even in the medical domain (Nori et al., 2023), we employed GPT-4o [^7] We use gpt-4o-2024-05-13 (Hurst et al., 2024). to perform a comparative evaluation of summaries generated by baselines and K-comp. All prompts utilized in the evaluation were structured using the identical format as presented in Table 12. Finally, we report examples of results for all baselines in Table 18 for qualitative analysis.
Query-Agnostic.
Figure 3 compares our approach to previous studies that compress prompts both with and without relying on the query. Intuitively, the baselines compared in query-agnostic section are compressed regardless of the question, which demonstrates that K-comp outperforms the other methods because it references the masked question. Detailed analyses are provided in Appendix B.
Figure 3: GPT-4o evaluation results with baselines. t-BioASQ indicates that MEDIQA was inferred using K-comp (or RECOMP) trained on BioASQ. For clarity, results in the 0% range are not indicated. Accordingly, we report all results in Table 17. (Above) Query-agnostic prompt compression, (Below) Query-based prompt compression methods, (Left) Results from the seen data, (Right) Results from the unseen data.
Query-Based.
SPLADE (Lassance and Clinchant, 2022) is a lexical-based retriever that reweights each term by emphasizing important terms associated with the query. The most informative top-1 passages among the top-5 retrieved passages are extracted and assumed to be compressed context. GPT-4o-mini 5 is the result of a summary generated using prompt 9, which is the same prompt used when synthesizing the RECOMP train data. The rationale for K-comp being rated higher than RECOMP is that even when trained on that dataset, it produces a summary that lacks specialization, which is similar to the output of FineTune in case study. By using a seq2seq model rather than token-level pruning such as LLMLingua, it allows the delivery of complete sentences to the reader LLM in QA tasks. This prevents the LLM from perceiving excessive noise, thereby ensuring relatively strong end-task performance among the baselines (Table 1). However, due to the maximum token limitation, the generated text is shorter and contains less information in the context, which results in lower evaluation scores for summary quality.
K-comp demonstrates comparable performance to GPT-4o-mini, particularly exhibiting significant superiority on MedQuAD. In query-based comparisons, GPT-4o-mini generates a rationale that is highly effective for answering based on its demonstrated capabilities in text generation. Similarly, K-comp exhibits performance comparable to that of GPT-4o-mini, despite being composed of only 2B parameters.
Unseen Evaluation.
In order to provide additional novelty to our approach, we evaluate baselines on data that was not used during training. Unlike query-based methods, such as RECOMP, which are trained to generate summaries optimized for responding to questions, our approach, based on the entities, demonstrates efficacy when applied to unseen data. The right-hand column of Figure 3 illustrates the results of applying each baseline to 140 test data from the MEDIQA (Ben Abacha et al., 2019). As discussed in Appendix B, Selective-Context (Li et al., 2023) and LLMLingua generate incomplete sentences, which result in inferior performance despite their compression in a question-agnostic fashion. In contrast, K-comp maintains a competitive performance on unseen data, comparable to the results evaluated on seen data.
A noteworthy point is the comparison result with query-based baselines. As can be seen in Table 8 and 9, the instructions used for synthesizing summaries in RECOMP and K-comp were as follows: "Compress … used to answer the question" and "Extract the content about the entity" respectively. By focusing on the entities, the objective of our training approach is to provide a concise context of the medical terminology that has been requested in the question. As a consequence, K-comp achieves a win rate that is similar to, and even exceeds, the results obtained in the training datasets. This presents a limitation of previous methods, which are unable to maintain their performance levels due to their reliance on the distribution of training data. On the other side, our training approach focuses on medical jargon, which is advantageous because the medical terminology remains consistent even when the data changes. Therefore, our causal knowledge injection substantially contributes to improving performance in data-scarce, closed-domain settings.
NLG metric.
Llama-8B Llama-70B Mixtral GPT-4o MedAlpaca Meditron
MedQuAD
Without compressor
Top-1 doc 2.96 3.11 2.63 4.27 1.65 3.14
Top-5 docs 3.22 3.36 3.07 4.73 1.85 3.53
With compressor
RECOMP 3.29 3.80 3.51 3.92 1.48 3.68
LLMLingua 2.41 2.87 2.53 3.68 1.61 3.21
FineTune 3.28 3.66 3.27 4.15 1.44 3.68
K-comp 3.55 3.66 3.73 4.48 1.98 3.85
MASH-QA
Without compressor
Top-1 doc 3.07 3.65 3.24 4.63 1.64 3.55
Top-5 docs 3.33 3.58 3.46 4.74 1.79 3.43
With compressor
RECOMP 3.68 4.14 4.49 4.15 1.34 4.01
LLMLingua 2.97 3.18 3.77 3.99 1.88 3.47
FineTune 3.36 3.79 3.78 3.93 1.74 3.51
K-comp 4.23 4.62 4.06 4.52 1.99 4.53
BioASQ
Without compressor
Top-1 doc 3.44 3.75 3.64 4.33 2.09 3.76
Top-5 docs 3.58 3.66 3.68 4.53 1.32 3.64
With compressor
RECOMP 3.64 3.92 4.10 3.58 1.18 4.00
LLMLingua 2.54 3.63 3.54 3.37 1.87 3.42
FineTune 3.63 4.05 3.59 4.01 1.48 3.88
K-comp 4.10 4.14 3.81 4.20 2.31 4.38
Table 5: Results evaluated using the G-Eval-4 metric (Liu et al., 2023b). The prompt used is presented in Table 13
The reliability of the methodology has been established through rigorous validation using the BertScore and UniEval metrics. To further reinforce its consistency, we employ G-Eval (Liu et al., 2023b), a state-of-the-art evaluation metric, to conduct a comprehensive supplementary assessment using a robust commercial LLM. In the case of G-Eval-4, GPT-4 is sampled 20 times, with the resulting average score used to minimize the impact of any potential variability. Given the extensive evaluation of our methodology across diverse datasets and a broad spectrum of models, conducting a full-scale assessment with G-Eval would entail substantial computational costs. Consequently, 1k data points per dataset were randomly selected as the test data. The results are reported in Table 5. Regarding performance, K-comp consistently exhibits superior performance in comparison to other baselines.
An interesting observation is that our results demonstrate a trend similar to that of UniEval. According to the G-Eval study, UniEval exhibited a higher correlation with human performance in summary, dialogue generation, and consistency evaluation than all baselines except G-Eval-4 (including G-Eval-3.5 with GPT-3.5 on occasion). This trend is also reflected in our findings. In Table 1, when the reader model is GPT-4o, it is evident that the answers generated by augmenting the top-5 documents have the highest scores, a phenomenon that is also observed in G-Eval. Furthermore, in the MedQuAD dataset, when the Llama-3-70B model is used as the reader, Uni-Eval exhibits a preference for RECOMP, which aligns with the corresponding G-Eval results. Similarly, in other cases, both metrics indicate a preference for K-comp. These findings suggest that UniEval effectively mirrors G-Eval, despite being a relatively older metric. Therefore, although human evaluation is not conducted in our study, the evidence suggests that our methodology is likely to correlate strongly with human judgments.
7 Conclusion
In this paper, we have proposed a novel method to improve retrieval augmented QA by compressing retrieved documents focused on the questions. We have devised a comprehensive scheme for identifying medical entities and automatically generating prior knowledge. This is followed by the extension of training and inference methods, which enable the autoregressive generation of summaries that incorporate domain knowledge while considering the context causally. Furthermore, we proved that this approach is practical even when applied to unseen evaluation, which represents a novel contribution in closed domains where data is scarce.
Limitations
Our methodology is limited in scenarios where the NER tool is unable to automatically detect ambiguous keywords or entities that are absent from the questions. To mitigate these issues, expanding the retrieval corpus with additional text chunks can inject more knowledge into the compressor and learn domain-relevant entities, but this will drastically increase the cost of annotating the data and require enormous resources for retrieval to perform nearest-neighbor searches. Therefore, we consider extending these retrieval datastores an important task in RAG, and this can be extended in future work.
Additionally, our study mainly focuses on English biomedical QA, which limits generalization to other languages and domains. Current studies in closed domains face challenges due to the scarcity of datasets, posing a considerable obstacle to the broader implementation of our methodology. We believe that, among closed domains, the medical QA has relatively more data, and we have proven our methodology in this domain. However, in other specific domains, not only QA data but also retrieval corpora are yet to be established. Furthermore, data availability in languages other than English is even more limited. Nevertheless, we recognize that our methodology has significant potential for extension to other languages and domains and that such expansion is necessary to demonstrate the generalizability of our training approach. Accordingly, we regard the application of retrieval-augmented QA in closed domains as a critical area of investigation, so we intend to extend our research to encompass additional domains in the future.
Ethical Considerations
In our research, we employed publicly available datasets, including MedQuAD, MASH-QA, BioASQ and MEDIQA. When synthesizing ground-truth summaries, we ensure that no personally identifiable information is used and that all data are anonymized. Our methodology is still in its early stages and is not yet suitable for direct practical use in medical domains, where reliability and accuracy are of utmost importance. In particular, hallucination can critically affect patient care and clinical decision-making. Therefore, our methodology is considered to mitigate hallucination by emphasizing domain knowledge in biomedical QA research, rather than substituting professional medical judgment, thus posing no risk of harm.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH))
This work was supported by Smart HealthCare Program(www.kipot.or.kr) funded by the Korean National Police Agency(KNPA, Korea) [Project Name: Development of an Intelligent Big Data Integrated Platform for Police Officers’ Personalized Healthcare / Project Number: 220222M01]
Appendix A Experimental Details
A.1 Model Details
To demonstrate that our contribution works universally regardless of reader models, we used a range of large language models (LLMs) (AI@Meta, 2024; Jiang et al., 2024; Han et al., 2023; Chen et al., 2023), including a commercial model (Hurst et al., 2024), with varying parameters and purposes. All open-source models were implemented based on HuggingFace’s Transformers (Wolf et al., 2020), and due to hardware constraints, we utilized AWQ (Lin et al., 2024a) models for LLMs with a substantial number of parameters. The specific model details are as follows:
Model HuggingFace’s Repository
K-comp google/gemma-2b
Retriever nomic-ai/nomic-embed-text-v1.5
Llama-3 8B meta-llama/Meta-Llama-3-8B
Llama-3 70B TechxGenus/Meta-Llama-3-70B-AWQ
Mixtral-8x7B TheBloke/mixtral-8x7b-v0.1-AWQ
MedAlpaca 13B medalpaca/medalpaca-13b
Meditron 70B TheBloke/meditron-70B-AWQ
Selective-Context meta-llama/Llama-2-7b-hf
LLMLingua meta-llama/Llama-2-7b-hf
SPLADE naver/efficient-splade-V-large-query
naver/efficient-splade-V-large-doc
A.2 Dataset Details
The datasets employed in the main experiments are MedQuAD (Ben Abacha and Demner-Fushman, 2019), MASH-QA (Zhu et al., 2020), and BioASQ (Krithara et al., 2023). MedQuAD encompasses a wide range of question types related to biomedicine, such as diseases, drugs, and medical tests. MASH-QA is a dataset from the consumer health domain where answers need to be extracted from multiple, non-consecutive parts of a long document. BioASQ is a biomedical dataset derived from PubMed, designed to support a range of tasks, including question-answering, information retrieval, and summarization. We obtained each dataset from the official websites provided by the papers (e.g., GitHub). In the case of the MedQuAD dataset, since there is no test data available, we randomly split the dataset into train/validation/test sets with an 80/10/10 ratio to conduct our experiments. Additionally, we used the MEDIQA (Ben Abacha et al., 2019) for further comparison with the baseline in GPT-4o evaluation section. MEDIQA comprises three tasks: Natural Language Inference, Recognizing Question Entailment, and QA, but only the QA task was used.
A.3 Implementation Details
We fine-tuned Gemma-2B (Team et al., 2024) with our knowledge injection objective and also used the FineTune baseline (FineTune) with the same model. Table 16 shows the hyperparameter settings that were used for training and inference. K-comp and FineTune were trained with the same hyperparameters and selected their optimal checkpoints based on the performance of the development set. We also trained the model via the Transforming Reinforcement Learning (Hu et al., 2024). When training on MASH-QA, we used 4 NVIDIA A100 GPUs with 80GB memory, otherwise we used 2 A100-80GB GPUs. For inference, we conducted experiments on a single A100-80GB GPU with vLLM (Kwon et al., 2023) to accelerate and efficiently perform inference.
Llama-3-8B Llama-3-70B Mixtral-8x7B GPT-4o MedAlpaca-13B Meditron-70B
BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval BertScore UniEval
MedQuAD
RAPTOR 74.96 60.84 77.66 71.79 66.4 61.49 84.20 57.78 83.59 48.42 74.67 60.30
K-comp 83.79 64.05 82.06 62.91 80.64 63.07 85.49 63.16 85.75 49.39 78.88 58.95
MASH-QA
RAPTOR 79.61 60.62 84.64 66.20 73.37 62.95 83.31 58.60 80.19 50.27 80.49 64.83
K-comp 83.45 61.45 83.64 61.40 81.49 62.91 84.49 64.06 81.95 51.76 82.67 61.85
BioASQ
RAPTOR 82.63 60.62 85.57 70.32 83.39 66.10 87.16 56.64 83.50 46.40 85.44 65.48
K-comp 86.61 62.28 87.03 62.89 86.52 66.31 87.77 64.38 84.88 45.47 86.70 62.90
Table 6: RAPTOR results. The implementation follows the same approach as that utilized for Table 1.
Appendix B Analysis of Query-Agnostic Section
Figure 3 compares our approach to previous studies that compress prompts without relying on the query, as illustrated in the top-left panel. GPT-4o-mini is the result of summarizing only passages without questions. We referred to the prompts used when training RECOMP, which can be seen in Table 10. Selective-Context (SC) (Li et al., 2023) is a method for the removal of contexts with low self-information at the token level. SC was compressed with Llama-2-7B (Touvron et al., 2023) as that used for LLMLingua. Both methods compress prompts by considering token-level dependencies, so the resulting sentences are incomplete when decoding. As can be seen in Table 18, LLMLingua generates sequences in which words and symbols are merely listed. At the same time, SC produces fragments due to token-level pruning, which also results in incomplete sentences. Consequently, although embedding-based metrics such as BertScore (Zhang* et al., 2020) or lexical-based metrics like ROUGE (Lin, 2004) may exhibit robust performance, they are limited in qualitative evaluations. Moreover, the methodology of compressing passages without referencing the query is unsuitable for retrieval-augmented QA tasks. Thus, both GPT-4o-mini and K-comp outperform the aforementioned baselines. In addition, it is expected that K-comp, by referencing masked questions, provides more relevant and specialized context compared to GPT-4o-mini, which does not reference the questions.
MedQuAD MASH-QA BioASQ
RAPTOR 20.7% 30.1% 43.4%
K-comp 77.9% 69.6% 52.5%
Tie 1.4% 0.3% 4.1%
Table 7: GPT-4o evaluation results with the RAPTOR approach.
Appendix C Additional baseline
In Section 5, a comparison of various baselines for RAG methods based on prompt compression is presented. However, prior RAG approaches that do not utilize compression were not implemented. Although the primary focus of our study is not to exhaustively explore non-compression-based methods, we introduce additional state-of-the-art (SOTA) baselines to facilitate a more comprehensive evaluation of the relative performance of K-comp. Specifically, RAPTOR (Sarthi et al., 2024) has been adopted as a baseline, and experiments have been conducted under several predefined conditions. Firstly, RAPTOR requires a summarizing LLM because it recursively summarizes to construct a tree structure. Therefore, we utilized FineTune model to perform context summarization, thereby eliminating the penalty for training. Secondly, RAPTOR has a limit on corpus size due to constraints in our computational resources required for summarizing clustered chunks of the retrieval corpus. Given the massive scale of the MedCorp (Xiong et al., 2024) corpus, we have replaced it with a corpus consisting of the top 5 documents retrieved by retriever across the entire dataset. It is contended that these experimental conditions do not disadvantage RAPTOR. Rather, we believe that restricting the corpus to the top-5 retrieved documents may offer potential advantages by improving the relevance and quality of the retrieved information.
As shown in Table 6, K-comp has demonstrated its effectiveness in QA tasks utilizing both prompt compression and non-prompt compression approaches. Moreover, as evidenced in Table 3, K-comp further underscores its cost efficiency, particularly in comparison to non-prompt compression methods. Additionally, we have conducted the GPT-4o evaluation described in Section 5, and the results are reported in Table 7. The outcomes indicate that the passages retrieved by RAPTOR approach are more frequently preferred in BioASQ than other datasets. As illustrated in Table 1, the FineTune model is optimally tuned for BioASQ compared to other datasets. This indicates that both overarching concepts and detailed information within the passages are effectively integrated during the tree construction process.
Notwithstanding, K-comp demonstrates higher accuracy than the state-of-the-art retrieval-augmented QA baseline (non-prompt compression), even though RAPTOR has an advantage over the methods presented in that it uses a corpus consisting of only the Top-5 docs. This validates our hypothesis that knowledge injection enhances QA performance. However, RAPTOR was originally implemented using closed models (GPT family), which may be weak for our work (Gemma).
Appendix D Licenses
MedQuAD, MASH-QA, BioASQ are licensed under the CC-BY-4.0, Apache-2.0, CC-BY-2.5 licenses, respectively. Gemma-2b, Nomic-embed-text-v1.5, Meta-Llama-3, Mixtral-8x7b-v0.1-AWQ, GPT-4o, MedAlpaca-13b, Meditron-70B-AWQ, Bge-reranker-large are licensed under the Gemma, Apache-2.0, Llama 3 Community, Apache-2.0, OpenAI, Creative Commons, Llama 2 Community, and MIT.
Appendix E Examples of Used Prompts
Please extract the content about the entity in fewer than four sentences.
### Passage
Therapies in Aicardi-Goutières syndrome.
Aicardi-Goutières syndrome (AGS) is a genetically determined disorder, affecting most particularly the brain and the skin, characterized by the inappropriate induction of a type I interferon-mediated immune response. In most, but not all, cases the condition is severe, with a high associated morbidity and mortality …(skip)
Treatments in Aicardi-Goutières syndrome.
Comprehensive reviews of the clinical characteristics and pathogenesis of Aicardi-Goutières syndrome (AGS), particularly its contextualization within a putative type I interferonopathy framework, already exist. However, recent reports of attempts at treatment suggest that an assessment of the field from a therapeutic perspective is warranted at this time …(skip)
Novel and emerging treatments for Aicardi-Goutières syndrome.
<bIntroduction</b: Aicardi-Goutières syndrome (AGS) is the prototype of the type I interferonopathies, a new heterogeneous group of autoinflammatory disorders in which type I interferon plays a pivotal role. The disease usually manifests itself during infancy, primarily affecting the brain and the skin, and is characterized by cerebrospinal fluid chronic lymphocytosis and raised levels of interferon-alpha and by cardinal neuroradiological features: cerebral calcification, leukoencephalopathy and cerebral atrophy …(skip)
Aicardi–Goutières syndrome
At the moment there are no therapies specifically targeting the underlying cause of AGS. Current treatments address the symptoms, which can be varied both in scope and severity. Many patients benefit from tube-feeding. Drugs can be administered to help with seizures / epilepsy …(skip)
Aicardi–Goutières syndrome
Treatment
### Entity
[research, clinical trial, Disorder]
Table 8: Prompt for synthesizing summaries used in K-comp.
Compress the information in the retrieved documents into a 2-sentence summary that could be used to answer the question:
Question: what research (or clinical trials) is being done for Aicardi-Goutieres Syndrome Disorde ?
Retrieved documents: Therapies in Aicardi-Goutières syndrome.
Aicardi-Goutières syndrome (AGS) is a genetically determined disorder, affecting most particularly the brain and the skin, characterized by the inappropriate induction of a type I interferon-mediated immune response. In most, but not all, cases the condition is severe, with a high associated morbidity and mortality …(skip)
Treatments in Aicardi-Goutières syndrome.
Comprehensive reviews of the clinical characteristics and pathogenesis of Aicardi-Goutières syndrome (AGS), particularly its contextualization within a putative type I interferonopathy framework, already exist. However, recent reports of attempts at treatment suggest that an assessment of the field from a therapeutic perspective is warranted at this time …(skip)
Novel and emerging treatments for Aicardi-Goutières syndrome.
<bIntroduction</b: Aicardi-Goutières syndrome (AGS) is the prototype of the type I interferonopathies, a new heterogeneous group of autoinflammatory disorders in which type I interferon plays a pivotal role. The disease usually manifests itself during infancy, primarily affecting the brain and the skin, and is characterized by cerebrospinal fluid chronic lymphocytosis and raised levels of interferon-alpha and by cardinal neuroradiological features: cerebral calcification, leukoencephalopathy and cerebral atrophy …(skip)
Aicardi–Goutières syndrome
At the moment there are no therapies specifically targeting the underlying cause of AGS. Current treatments address the symptoms, which can be varied both in scope and severity. Many patients benefit from tube-feeding. Drugs can be administered to help with seizures / epilepsy …(skip)
Aicardi–Goutières syndrome
Treatment
Compressed documents:
Table 9: Prompt for synthesizing summaries used in RECOMP training and query-based GPT-4o-mini.
Compress the information in the retrieved documents into a 2-sentence summary
Retrieved documents: {{Top-5 retrieved passages}}
Compressed documents:
Table 10: Prompt for synthesizing summaries used in query-agnostic GPT-4o-mini.
### Passage
Aicardi-Goutières syndrome (AGS) is a genetically determined disorder affecting the brain and skin, characterized by inappropriate immune responses due to type I interferon. Current research focuses on understanding its pathogenesis and developing targeted therapies, with some recent attempts exploring treatments like Janus kinase inhibitors and anti-IFN-$\alpha$ antibodies. Despite advancements, there are still challenges in assessing efficacy and addressing open questions related to treatment effectiveness. Ongoing clinical trials aim to evaluate the efficacy of new therapies and address the underlying causes of AGS. Overall, ongoing research and clinical trials are essential for developing effective treatments for AGS and type I interferonopathies.
### Entity
research: Systematic study undertaken to increase knowledge
clinical trial: Phase of clinical research in medicine
Aicardi-Goutières Syndrome: Aicardi-Goutières syndrome is a disorder that mainly affects the brain, the immune system, and the skin.Most newborns with Aicardi-Goutières syndrome do not show any signs or symptoms of the disorder
### Questions
what research (or clinical trials) is being done for Aicardi-Goutieres Syndrome Disorde ?
### Passage
Therapies in Aicardi-Goutières syndrome.
Aicardi-Goutières syndrome (AGS) is a genetically determined disorder, affecting most particularly the brain and the skin, characterized by the inappropriate induction of a type I interferon-mediated immune response. In most, but not all, cases the condition is severe, with a high associated morbidity and mortality …(skip)
Treatments in Aicardi-Goutières syndrome.
Comprehensive reviews of the clinical characteristics and pathogenesis of Aicardi-Goutières syndrome (AGS), particularly its contextualization within a putative type I interferonopathy framework, already exist. However, recent reports of attempts at treatment suggest that an assessment of the field from a therapeutic perspective is warranted at this time …(skip)
Novel and emerging treatments for Aicardi-Goutières syndrome.
<bIntroduction</b: Aicardi-Goutières syndrome (AGS) is the prototype of the type I interferonopathies, a new heterogeneous group of autoinflammatory disorders in which type I interferon plays a pivotal role. The disease usually manifests itself during infancy, primarily affecting the brain and the skin, and is characterized by cerebrospinal fluid chronic lymphocytosis and raised levels of interferon-alpha and by cardinal neuroradiological features: cerebral calcification, leukoencephalopathy and cerebral atrophy …(skip)
Aicardi–Goutières syndrome
At the moment there are no therapies specifically targeting the underlying cause of AGS. Current treatments address the symptoms, which can be varied both in scope and severity. Many patients benefit from tube-feeding. Drugs can be administered to help with seizures / epilepsy …(skip)
Aicardi–Goutières syndrome
Treatment
### Questions
what research (or clinical trials) is being done for Aicardi-Goutieres Syndrome Disorde ?
Table 11: Prompt for reader LLMs. (Above: K-comp, Below: Top-5 passages)
Select the summary (Summary 1, Summary 2, Summary 3, or Summary 4) that is more relevant and informative as a rationale for the given question.
In particular, biomedical QA requires expertise to be credible, so choose a summary where expertise exists in the domain.
Choice: [Summary 1, Summary 2, Summary 3, Summary 4, Tie], Do not offer any opinions other than the choice.
### Summary 1
{{summary 1}}
### Summary 2
{{summary 2}}
### Summary 3
{{summary 3}}
### Summary 4
{{summary 4}}
### Question
{{question}}
Table 12: Prompt for GPT-4o evaluation. The order of summaries was randomized throughout the evaluation process to mitigate potential bias associated with their positioning.
You will be given a question, passage, and answer (both the provided answer and the correct answer).
Your task is to evaluate how coherent and relevant the provided answer is based on the question, passage, and correct answer. The passage is provided as a reference to determine if the provided answer appropriately aligns with the information required to answer the question.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
Evaluation Criteria:
Coherence (1-5) - the degree to which the provided answer is logically consistent, well-structured, and directly relevant to the question. The answer should make logical sense and align with the correct answer, while addressing the question clearly and effectively.
Evaluation Steps:
1. Read the question carefully to identify the specific information it seeks.
2. Read the answer and compare it to the correct answer. Check if the provided answer addresses the question directly, aligns logically with the passage’s content, and is presented in a clear, organized, and complete manner.
3. Assign a score for coherence on a scale of 1 to 5, where 1 is the lowest and 5 is the highest based on the Evaluation Criteria.
Example:
Passage:
{{passage}}
Question:
{{question}}
Correct Answer:
{{correct answer}}
Answer:
{{answer}}
Evaluation Form (scores ONLY):
Coherence:
Table 13: Prompt for G-Eval-4 evaluation metric (Liu et al., 2023b).
Appendix F Extended Case Study
K-comp
### Entity
x-ray: Form of short-wavelength electromagnetic radiation
rheumatoid arthritis: Type of autoimmune arthritis
### Summary
X-rays are the most common method for assessing the degree of joint destruction in rheumatoid arthritis (RA), revealing osteopenia and joint space narrowing as the disease progresses. They are particularly useful for identifying bony erosions and assessing treatment efficacy. While imaging techniques like MRI and ultrasound are also employed, X-rays remain the most common for evaluating joint damage in RA. Early diagnosis and treatment are crucial to modify the disease’s course and reduce complications. Understanding the disease’s pathophysiology and imaging findings is essential for effective diagnosis and management.
Llama-3-8B’s Answer
X-rays diagnose rheumatoid arthritis (RA) by revealing changes in the axial skeleton and joint space. As the disease progresses, early RA may not show significant changes on X-rays, but it can indicate osteopenia and joint space narrowing. Advances in imaging techniques like MRI and ultrasound enhance the ability to detect soft tissue changes and joint damage, which is crucial for accurate diagnosis and treatment management. Overall, X-rays serve as an important tool for evaluating the structural changes associated with RA.
Llama-3-70B’s Answer
X-rays can reveal joint space narrowing and osteopenia, which are common findings in RA.
X-rays can show early signs of joint damage in RA, such as bone erosion and synovial inflammation.
X-rays can detect soft tissue changes and joint damage in RA, which is important for diagnosis and treatment management.
X-rays can assess the degree of joint destruction in RA, including changes in the axial skeleton and joint space.
FineTune
### Summary
Rheumatoid arthritis (RA) is a chronic systemic disease characterized by inflammation of the synovial tissue, affecting 0.5%-1.0% of the global population, with a higher prevalence in females. Modern diagnostic techniques, including X-rays, MRI, and ultrasound, are crucial for assessing joint damage and evaluating treatment efficacy. Early diagnosis and treatment are essential to modify the disease’s course and reduce severe complications. Advances in imaging techniques, particularly MRI, enhance the ability to detect and manage RA more effectively. Understanding the disease’s pathophysiology and various imaging findings is essential for accurate diagnosis and tailored treatment strategies. Overall, imaging plays a significant role in the management and monitoring of RA.
Llama-3-8B’s Answer
X-rays are used to diagnose RA by detecting joint damage and erosion.
Llama-3-70B’s Answer
X-rays are a valuable tool in the diagnosis of RA, as they can detect joint damage and erosions caused by the disease. In the early stages of RA, X-rays may not show any abnormalities, but as the disease progresses, they can reveal joint space narrowing, erosions, and deformities. X-rays can also help monitor the progression of the disease and assess the effectiveness of treatment. However, X-rays have limitations, as they may not detect early changes in the joints or soft tissue inflammation. Therefore, other imaging modalities, such as MRI and ultrasound, may be necessary to provide a more comprehensive assessment of RA.
Table 14: Case study in section 4. (Above: K-comp, Below: FineTune)
Appendix G Dataset Statistics
MedQuAD MASH-QA BioASQ
Train Validation Test Train Validation Test Train Validation Test
Original 13,127 1,640 1,640 27,728 3,587 3,493 3,209 803 707
After filtering 9,064 1,095 1,554 20,458 2,637 3,226 2,284 563 647
% Filtered 30.9 33.2 5.2 26.2 26.4 7.6 28.8 29.8 8.4
Table 15: Dataset sizes before and after filtering in the entity recognition step. For test data, filtering is applied exclusively to questions lacking any entities. For other datasets, filtering is additionally conducted for the absence of corresponding descriptions for the recognized entities.
Appendix H Hyperparameters
Hyperparameters MedQuAD MASH-QA BioASQ
Global batch size 8
Gradient accumulate 1
Learning scheduler cosine
Weight decay 3%
Epochs 3
Temperature 0.01
Top_p 1.0
Early stopping True (valid loss)
Optimizer AdamW (Loshchilov and Hutter, 2019)
($\beta_{1}$=0.9, $\beta_{2}$=0.999, $\epsilon$=1e-8)
Learning rate 1e-4 1e-4 1e-5
Table 16: Hyperparameters used in the experiments.
Appendix I GPT-4o Evaluation Results
Method Seen Data Unseen Data
MedQuAD MASH-QA BioASQ
MEDIQA
(t-MedQuAD)
MEDIQA
(t-MASH-QA)
MEDIQA
(t-BioASQ)
Query-Agnostic Selective-Context 0 0 3 0 0 0
LLMLingua 4 10 1 0 0 1
GPT-4o-mini 259 635 202 25 19 48
K-comp 1290 2581 441 115 121 91
Tie 1 0 0 0 0 0
Query-Based SPLADE 88 371 70 16 17 19
RECOMP 252 406 82 6 8 4
GPT-4o-mini 504 1240 253 56 51 71
K-comp 710 1208 242 62 64 46
Tie 0 1 0 0 0 0
Table 17: All results of the GPT-4o evaluation.
Appendix J Examples of Compressed Context Generated by Baselines
Question: Are there any adults who shouldn’t get the meningococcal vaccine?
Selective-Context
Meningococcal disease Routine primary immunization against who planning Although conjugate vaccines in Since safety to Prevention Invasive disease The disease can rapidly or despite with required household Although cases often Serogroup B accounts but contained Adolescents disease Because in contained they potentially recently containing This T-cell-dependent vaccine that should The Advisory Committee to entering The vaccine also may belong Meningococcal disease through caused While IMD incidence peaks both infants carriage rates often increasing Effective vaccines Because represent cases often have the highest carriage rate efforts in This review describes and characteristics With countries that will also Rates increasing. The burden due available older. Incorporation into offers affect The routine use will Awareness attitudes among among Meningococcal- under However MenB vaccination coverage We investigated MenB disease among An online survey parents among There were 606 total participants parents of MenB (<in</i = also Non-vaccinated cohorts reported ( though However all cohorts reported high interest MenB disease Vaccinated cohorts relied had SCDM at though those Barriers included PCP recommendation uncertainty There gaps MenB disease SCDM patients resulting discussing Targeted education on MenB may these opportunities
LLMLingua
Men ine most or are. Althoughss conjug. Since established disease In be ande.otic is for. population. B the.s Becauseoccaloups are the, are The This and The to who to certain the isidence in, are,., haveination, regardingterm of in ofents. on with have ofis The limitations been populations In more TheA of and3old under population regardingold of (<ini andination, and SCDM among parents and patients in the US, resulting in missed opportunities for discussing and administering MenB vaccination. Targeted education on MenB and vaccination recommendations may increase these opportunities and improve MenB vaccination awareness and initiation.
GPT-4o-mini (query-agnostic)
Routine primary immunization against meningococcal disease is advised for adults in endemic areas or traveling there, with conjugate vaccines preferred for those under 55, while polysaccharide vaccines are used for older adults due to safety concerns. Adolescents and young adults are prioritized for vaccination due to their higher incidence and carriage rates of the disease, yet awareness and vaccination coverage for serogroup B remains low, necessitating targeted education to improve understanding and uptake.
SPLADE
Meningococcal disease Routine primary immunization against meningococcal disease is recommended for most adults living in areas where meningococcal disease is endemic or who are planning to travel to such areas. Although conjugate vaccines are the preferred meningococcal vaccine in adults 55 years of age or younger, polysaccharide vaccines are an acceptable alternative for adults in this age group if the conjugated vaccine is unavailable. Since safety and efficacy of conjugate vaccines in adults older than 55 years of age have not been established to date, polysaccharide vaccines should be used for primary immunization in this group.
RECOMP
Adults who should not receive the meningococcal vaccine include those who are 55 years or older, those in high-risk groups, and those who are pregnant or breastfeeding. Additionally, polysaccharide vaccines are recommended for adults over 55, as conjugate vaccines are not recommended for this age group due to safety concerns.
GPT-4o-mini (query-based)
Most adults are recommended to receive the meningococcal vaccine, particularly those living in endemic areas or traveling to such regions, but adults over 55 years should use polysaccharide vaccines due to the unestablished safety and efficacy of conjugate vaccines for this age group. Overall, while there are no strict categories of adults who should not get the vaccine, individuals with specific health concerns or those who have experienced adverse reactions to prior vaccinations should consult with healthcare professionals for personalized guidance.
K-comp
### Entity
people: Plurality of persons considered as a whole
meningococcal vaccine: Vaccine against meningococcal disease
### Summary
The meningococcal vaccine is recommended for adults in endemic areas and for adolescents and young adults in high-risk groups. The meningococcal vaccine is safe and effective, but the polysaccharide vaccine is preferred for those 55 years and older due to safety concerns in older adults. Vaccination efforts are crucial, as awareness and uptake of the vaccine remain low among adolescents and young adults, particularly in the United States. Education and targeted efforts are needed to improve awareness and vaccination rates among this population. The vaccine is particularly important for preventing invasive meningococcal disease, which can have severe consequences. Overall, effective vaccination strategies are essential to reduce the incidence of meningococcal disease.
Table 18: Examples of summaries generated by all the baselines